A solution consists of 20% acid, and the remainder is water. A chemist wishes to strengthen this solution. How many quarts of acid must be added to 24 quarts of the solution to make a mixture containing 40% acid? show solution
Answer by richwmiller(17219)

We have 24 qt of 20% acid
We want to add some 100% acid to make 40% acid
1*x+0.2*24=0.4(24+x)
1*x+4.8=9.6+0.4x
1x-0.4x=9.6-4.8
0.6x=4.8
x=8 qt at 100% (8) is added to 24 qt at 20%(4.8) to make 40% of 32 qt(12.8)
check
1*8+0.2*24=0.4(24+8)
8+4.8=0.4(32)
12.8=12.8
ok